OpenStudy (anonymous):

Which number can replace the variable to make these fractions equivalent? n/9=56/63 = A. 2 B. 7 C. 8 D. 9

OpenStudy (anonymous):

medal and fan :)

OpenStudy (anonymous):

c

OpenStudy (anonymous):

can you help me with 2 more

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which is a correct comparison of the fractions? 4/5and5/6 A. B. C.

OpenStudy (anonymous):

you still there

OpenStudy (anonymous):

what are the options

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

A.4/5<5/6 B.4/5=5/6 C.4/5>5/6

OpenStudy (anonymous):

k

OpenStudy (anonymous):

one more

OpenStudy (anonymous):

4/5=.8 5/6=.83 so a

OpenStudy (anonymous):

Which pair of fractions shows 5/6 and3/8 correctly converted with their least common denominator? A.5/48and3/48 B.40/18and18/48 C.5/14and3/24 D.20/24and9/24

OpenStudy (anonymous):

lcd=24 D

OpenStudy (anonymous):

hold on im submitting it

TheSmartOne (thesmartone):

@koook This is my 5th time telling you That this is against the CoC. You are giving direct answers away. Please read the CoC before mods suspend you for this. Thank you. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

i do k12

OpenStudy (anonymous):

it was

TheSmartOne (thesmartone):

It was a test?

OpenStudy (anonymous):

i got it though

OpenStudy (anonymous):

pretest not importent test

OpenStudy (anonymous):

optional

OpenStudy (anonymous):

i got a 5/5 thanks kook

OpenStudy (anonymous):

no problem. happy to help :)

OpenStudy (anonymous):

heres your medal

TheSmartOne (thesmartone):

You know that posting even pretest questions is not allowed right? It says it in the terms and condition. Thank you. Also be careful there are mods from Online schools making sure that no one cheats. http://openstudy.com/terms-and-conditions

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

well this is an assessment does that count

TheSmartOne (thesmartone):

Well cheating in general is not allowed. You were suppose to ask a question and cooperate with koook and solve the problem. Then you could post your second question and ask koook if you solved it correctly and from there on checked your answers. Not getting the answers.

OpenStudy (anonymous):

im about to shut this question down :/ sorry this is my first time and i normally dont read terms of conditions

TheSmartOne (thesmartone):

Thank you and have a great day.